# Modeling when and how physics PhD students search for a research group: the role of interests and prior research experiences in timely group integration

### v2

@@ IV.5 Difficulties and concerns in the group search
Another highly cited concern for many students ($N=22$) was Doubting ability to do research. These students described feeling unprepared for graduate research, and often referred to a sense of imposter syndrome, which is a common phenomenon amongst PhD students [86, 87]. As Wendy put it, “Just a difficulty, maybe it’s more of like imposter syndrome. Like, let’s say, I do join the research group, am I going to do well? Am I going to be smart enough to handle all these smart and complex topics?” Meanwhile, when Pauline changed to a new research group at the end of her second semester, she said, “I can’t say I felt as confident because it was all, it’s a lot of in-lab work,” which she had never done before. This code also included instances of students discussing difficulties understanding what a group’s research entailed, which we detailed in our previous work [31].
@@
These concerns were less common among students who joined a group earlier, who were also more likely to be more confident in the research they wanted to pursue. One major contributor to students’ self-efficacy is experience successfully performing a task (e.g., a student who does research in an optics lab tends to feel more capable of doing that research in the future) [43]. Students who enter graduate school having already done a significant amount research of research is more likely to have more strongly developed interests, and is also more likely to feel confident in their ability to succeed in future research.
@@
Some students ($N=16$) also described Difficulty communicating with advisors as an issue they experienced during their search for a group. This difficulty predominantly revolved around students’ own anxiety being a barrier to reaching out to prospective advisors. Eric noted that worrying about impressing professors “wasn’t a huge barrier” but “definitely made me hesitate in sending emails.” Meanwhile, Dev lamented that the process of contacting advisors “feels asymmetric” since “we have to put so much effort to figure out what they’re doing, while some of them don’t even update their websites… there’s no way to communicate except for them feeling like replying to our emails.” Students who committed to a group first semester or earlier encountered this difficulty at a slightly lower rate. However, none of these students shared the concern Wanting more ways to meet grad students and groups groups, whereas six of the students who joined second semester or later were found under both codes. This may suggest that the subset of students who joined later in the first year were more likely to put off reaching out, and therefore tended to want departmental support with that process.
@@
Indeed, we see that students who joined second semester or later were far more likely to discuss Wanting more ways to meet grad students and groups. Students who sought more structure meeting groups often talked about wanting ways to try different research, which aligns with the finding that these students also tended to enter graduate school with broader research interests and were less likely to have already connected with individual research groups. Additionally, we see that this subset of students also tended to have more struggles with Coursework and research in tension, as classes prevented them from dedicating time to their group search. Students who joined early occasionally cited this difficulty, but it was typically in regard to their desire to begin doing research in their lab rather than taking classes that they felt were not beneficial. Students who were able to enter into a research group early did not have to search for a group while taking classes. This lack of time may have played a role in wanting more structured ways to meet research groups, as their schedules prevented them from doing so on their own time.
@@
Although many of the difficulties discussed more commonly affected students who had not joined a group by their first semester, our results also reveal several drawbacks of joining a research group so early (the bottom of Table IV.5). Most notably, there was a higher prevalence of students being worried about Commitment to research before joining a group. Olivia, who planned to begin doing research with a professor as soon as she arrived in graduate school, feared that she could “get there and start talking about the research and doing it, and I might not like it at all… I don’t really know the solution for that.” Oftentimes students who had committed early to a group worried that they may be missing out on a better opportunity. This included students who expressed being confident in their choice of research. Ursula recalled that she “was trying not to pigeonhole myself too early still, even though I knew I loved this subfield.” In the end though, she committed to working with a professor whom she collaborated with during her undergraduate research.
@@
IV.5.1 Looking at the subset thinking about leaving
@@
As the first year progressed, both students experienced difficulties with Funding, Wanting more ways to meet grad students and groups, Coursework and research in tension, Difficulty communicating with advisors, and that contributed to their consideration to leave. Being denied from groups due to funding constraints negatively impacted sense of belonging for both students. Carmen was rejected from the first group they approached: “being my first attempt to reach out to a potential advisor, that was not a good feeling.” Brianna went to four professors in her department who all“said no, because they didn’t have funding for another grad student… personally, I thought I was cursed.” Carmen felt that the department did not do an adequate job facilitating interactions between graduate students, saying “I also felt like I lacked a sense of community… yeah, a lot of my concerns were community based.” Brianna recalled being confused as to “whether or not I had to find [a group] right away, as opposed to like, waiting out and focusing on trying to pass the courses first.” Opting to focus on coursework, she “was under the impression that could wait till my second year,” until the department informed her she needed a group by the summer, which made her feel“taken aback. I was scared.” Carmen described all of the first year responsibilities as “more overwhelming than I had expected.” Both described anxiety associated with seeking out help from faculty; as Brianna said, “I just didn’t know what was appropriate to ask and what not to ask… when you’re talking to people who hold the fate of your career in their hands, it’s like, I don’t know if they’re gonna think I’m dumb for asking this question. Because I’ve never experienced any of this before.”
@@
Leading up to graduate school, Tabitha and Kwan described antipodal experiences to those described by Brianna and Carmen. Both Tabitha and Kwan came into graduate school knowing exactly the kind of research they wanted to pursue, and had been in contact with prospective advisors before arriving to their programs. Kwan was guided to potential groups by his undergraduate research advisor, saying “some other candidate professors made my list but I screened them out by talking with my previous advisor. He was very supportive for applying to graduate school.” Meanwhile, Tabitha’s prospective advisor was a co-author on one of her undergraduate research papers: “So one of the professors at [this school] reached out to me and he was like, ‘Hey, I read your thesis. We would like to keep on working on this and make it an actual paper’ and I was like, sure.” Thus, when it came time for her to find a PhD program, this faculty member was a natural fit. She recalled him asking, “‘Have you to considered going to [this school]?’ Because he just happened to work here. Once again it’s like, you want me, you want to have me as your student? I’m not gonna say no.”
@@
Tabitha’s and Kwan’s research topics matched their interests excellently, but their work environments drove them to leave. From the outset Kwan felt that it was “inappropriate” for his advisor’s significant other to be working in the lab, and felt that the advisor treated him unfairly because of this dynamic. “It was a kind of big problem, a big burden. And one day we just yelled at each other, this kind of serious argument.” In his next meeting, his advisor said that their research style was “not fit” for him, and asked him to leave the lab. Kwan, an international student, recalled, “it was really stressful… I thought, why did I come to US?” He continued, saying “…it was a kind very tough time for me. And I regretted about my decision to come to US a lot during my first year.” Meanwhile, Tabitha became disillusioned with her group due to the poor working environment fostered by the advisor. She said her advisor “created a very toxic work environment… always trying to gaslight me into believing I was not good enough to be here… [trying] to force me to work outside of my working hours.” Yet because she had come to the school specifically to work with this professor, she felt obligated to stay: “It took me a while to realize that I was not in a healthy work environment. Mostly because I was so, in a way, indebted to this professor for giving me the opportunity to come here. I think I thought, ‘Oh, yeah. He’s not great, but I’m here because of him.” More details of Tabitha’s case are detailed in our prior work [31].
@@
Other students in our study had joined groups early and subsequently switched, but only Tabitha and Kwan indicated that they considered whether they would be able to continue in their programs. Other students in our sample were motivated to switch groups in order to find research that better aligned with their interests, but none of them indicated that their experience drove them to consider leaving their programs. Indeed, for students who came in with broad interests and joined a group in order to try out the research, switching was an expected and positive move. Consistent with prior literature on the impact of negative advising relationships, Tabitha’s and Kwan’s narratives demonstrate the major impact of group culture on student satisfaction. They also illustrate the negative influence that joining a group so early can have on students’ likelihood to stay in their programs. Tabitha said of her experience, “I definitely think that could have been avoided,” and she wished that that someone had told her earlier to rethink working for her advisor. Reflecting on her search for a group, she emphasized to future students that “It’s a balance. It’s hard to balance having a good advisor and having a research field that you’re actually passionate about.”
@@
Figure 6: A model of characteristic pathways that physics PhD students take during the process of evaluating the congruence of a group’s research interests, exposing several underlying factors that contribute to students to thinking about leaving their programs. Whereas prior research has correlated individual constructs (e.g., broad interests) with leaving [1], this model illustrates how these factors interact and influence a range of student outcomes. Although it was not the focus of the current study (signified by the light grey outline), we include evaluating group culture as a known feature of the group search process. See Figure 7 for examples of how students in this study map onto the model.
Figure 7: Several examples of students whose experiences searching for a group with congruent research interests are well-described by the model in Figure 6. Each row represents a characteristic type of student experience, described alongside the student names in the study who fit that archetype. The paths listed under a) and b) describe students who our results suggested were at a higher risk of leaving their programs.
@@ V Discussion
This study has examined the ways that physics PhD students’ research interests influence their search for a research group in graduate school. In this section, we synthesize our results in an explanatory model of student outcomes, shown in Figure 6. Grounded in data, the model highlights the connections between students’ backgrounds, their research interests, when they search for a group, and the consequences of searching for a group before or during graduate school. Moreover, by tracing students’ journeys group back to their undergraduate years, this model offers unique insight into some of the ways that inequitable access to formative experiences for can affect outcomes down the road in graduate school. Whereas prior research has correlated individual constructs (e.g., broad interests) with leaving [1], this model illustrates how these factors interact and influence a range of student outcomes. Figure 7 provides several examples of how students described throughout the study map onto the diagram shown in Figure 6. The elements of Figure 7 align from left to right with the elements of the path diagram, and share headings to highlight how the two figures relate.
@@
Critically, our work offers insight into two characteristic pathways that we observed as placing students at a higher likelihood of leaving their programs, highlighted in Figs.7a) and b). Figure 7a) represents the path exemplified by Brianna and Carmen. Both were first-generation students who entered graduate school unsure of what research they wanted to pursue. Thus their group search took place in graduate school, where they encountered a number of difficulties that impacted their overall sense of belonging in the program. Indeed, both considered leaving their programs before finding a research group. Meanwhile, Figure 7b) represents the path exemplified by Tabitha and Kwan, beginning with their strong research backgrounds that made them confident in the research they wanted to do in graduate school. Both students found a research advisor before they arrived in graduate school, primarily based on their research topic. However, upon working in their research groups during their first year, they recognized that their labs did not provide a working environment conducive to their success and considered leaving.
@@
Many other students’ experiences are described well by the model too, as summarized in Figure 7. For example, Elena, Nina, and Alex are illustrative of the path shown in Figure 7c). Their undergraduate research played a defining role in determining the research they wanted to pursue in graduate school. Moreover, although this analysis primarily focused on the role of research interests in students’ search for a research group, many students in our sample were clearly cognizant of cultural factors in their choice of research group. For Elena, Nina, and Alex, their undergraduate institutions and research experiences gave them strong connections to their prospective groups. Ursula and Elena had already worked on the same collaborations as their eventual graduate advisors, and Alex’s undergraduate advisor had previously worked with his prospective graduate advisor. Thus, this subset of students had insight into both the research and cultural aspects of their research groups early on in their search process, and they joined their labs with high confidence in having a good fit.
@@
Another notable pathway through the model is illustrated by Eric, Cole, and Kali in Figure 7d). We referenced Eric and Cole in Section IV.2 as having a broad set of interests coming into graduate school, but who joined a research group early anyway. Kali was also discussed in this section, as she indicated that she wanted to “explore.” Thus, none of these students applied to their schools with the intention of working with a specific professor. Eric and Cole joined groups near the beginning of their graduate programs anyway, figuring that if it did not work out they would have time to switch. In effect, they were creating their own “rotation systems” Meanwhile, Kali’s program had institutionalized a rotation program, which also allowed her to try a lab and switch. We did not include Kali in our count of students who switched groups since it was a mandatory aspect of her graduate program, but we include her experience here since it most closely resembles the path shown in Figure 7d). Indeed, all three students found their first labs to be poor fits, predominantly due to the research. Therefore, they left and found new groups by the end of their first year summer.
@@
Lastly, in Figure 7e) we highlight several students who took a similar path as Brianna and Carmen but who never reported feeling isolation in their programs. Rather, they were able to navigate the search process in spite of the difficulties they experienced. Isaac and Ash had both done undergraduate research but remained unsure of what exactly they wanted to do in graduate school. Both reported having to balance coursework and research was a major issue during their first year and hindered their search, and both indicated that their departments did not provide much support in helping facilitate their group search. However, they never indicated that they were considering leaving their programs, and navigated their difficulties to find research groups by the end of the first year. One important factor that undoubtedly contributed to this difference was Brianna and Carmen’s status as first-generation students. They cited this several times as a major reason for their struggle adjusting to graduate school overall. Difficulties navigating other aspects of graduate school while also worrying about finding a group undoubtedly exacerbated their negative feelings. We suspect that other factors contributed to the different outcomes as well, including departmental culture and the presence or absence of a strong graduate student cohort, but we did not investigate these elements in this work. In the future, we hope to examine differences across students’ narratives to investigate what students found most helpful and hurtful while looking for a group.
@@
Meanwhile, handbooks often have clear expectations for passing classes and qualifying exams during the first year. It is therefore unsurprising that students who come in relying on formal departmental guidance to navigate the first year may deliberately opt to delay looking for a group in favor of passing classes. However, this systematically disadvantages the 25% of physics graduate students who identify as first-generation college students, as this group is less likely to reach out for help on their own [89, 90]. Lacking guidance to the contrary, students may perceive experiences such as being denied access to a group due to funding seem like personal failings, contributing to feelings of isolation.
@@
Limitations and future research Interviewing students during their first and second years of physics graduate school allowed us to obtain deeper insights into their thoughts “in the moment” while looking for a group. However, in doing so we sacrificed some insight into these students’ outcomes later on in their graduate careers. The grey arrows in Figure 6 represent areas of the model that we did not investigate in this paper, but that we suspect to exist. One area that we did not have enough data to investigate is the outcomes for students who join a research group later on but experience a poor fit. Realizing that a lab is a poor fit can take time, as exemplified by Tabitha and Kwan who worked in their labs for over a year before switching. Students who join a group later in their first year might not realize that the fit is not great until late into their second year. By then, the cost of switching might feel too great, and could leave them feeling as though they have no choice but to stay in the poor research environment or leave the program. More research into these students’ experiences would be needed to understand how they fit into our model.
@@
The goal of this paper from the outset was to emphasize the role of students’ research interests in how they find a group. Thus, a clear limitation of the results presented here is that they do not offer as much insight into how students evaluate group cultures and working environments. A positive interpersonal dynamic in a research group is clearly essential for students to thrive, but how students weigh these considerations in comparison to their research interests remains unexplored within our data. Previous work in the context of biology graduate education indicates that students weighed both factors heavily when choosing a group [27], and we see indications of that in our data as well. We hypothesized how this step might fit into our current model with a grey outlined box in Figure 6. However, a fuller description is needed to understand how students approach finding a group with a positive culture.
@@
Lastly, our data does not include students who left their graduate programs. Their perspectives on the physics graduate experience are critically important, but are often difficult to collect due to poor tracking of students once they leave their programs [1]. Although our data included students who reported a strong consideration of leaving, they ultimately stayed. Gaining a more accurate characterization of how finding a research group impacts student retention, satisfaction, and leaving will require more work to gather data from those who did not complete the PhD. Future research should also attend to how it is experienced across demographic groups and within different institutional contexts, as our study was limited in these regards. Our sample was diverse in many ways, but our protocol did not probe how students felt their identities impacted their group search. Moreover, this analysis did not systematically examine differences across race, gender, or international status regarding time to join groups and development of research interests. And although first-generation status came up naturally during our study played a significant role in several students’ narratives, this was not our central focus. Future research must attend to how physics students’ identities shape the first year of the doctoral experience in order to provide more holistic and equitable support for all students.

### v3

@@ IV.5 Difficulties and concerns in the group search
Another highly cited concern for many students ($N=22$) was Doubting ability to do research. These students described feeling unprepared for graduate research, and often referred to a sense of imposter syndrome, which is a common phenomenon amongst PhD students [90, 91]. As Wendy put it, “Just a difficulty, maybe it’s more of like imposter syndrome. Like, let’s say I do join the research group, am I going to do well? Am I going to be smart enough to handle all these smart and complex topics?” Meanwhile, when Pauline changed to a new research group at the end of her second semester, she said, “I can’t say I felt as confident because it was all, it’s a lot of in-lab work,” which she had never done before. This code also included instances of students discussing difficulties understanding what a group’s research entailed, which we detailed in our previous work [32].
@@
These concerns were less common among students who joined a group earlier, who were more likely to have higher confidence in the research they wanted to pursue. One major contributor to students’ self-efficacy is experience successfully performing a task (e.g., a student who does research in an optics lab tends to feel more capable of doing that research in the future) [45]. Students who enter graduate school having already done a significant amount research of research are more likely to have more strongly developed interests, and are also more likely to feel confident in their ability to succeed in future research.
@@
Some students ($N=16$) also described Difficulty communicating with advisors as an issue they experienced during their search for a group. This difficulty predominantly revolved around students’ own anxiety being a barrier to reaching out to prospective advisors. Eric noted that worrying about impressing professors “wasn’t a huge barrier” but “definitely made me hesitate in sending emails.” Meanwhile, Dev lamented that the process of contacting advisors “feels asymmetric” since “we have to put so much effort to figure out what they’re doing, while some of them don’t even update their websites… there’s no way to communicate except for them feeling like replying to our emails.” Students who committed to a group first semester or earlier encountered this difficulty at a slightly lower rate. However, none of these students shared the concern Wanting more ways to meet grad students and groups, whereas six of the students who joined second semester or later were found under both codes. This may suggest that the subset of students who joined later in the first year were more likely to put off reaching out, and therefore tended to want departmental support with that process.
@@
Indeed, we see that students who joined second semester or later were far more likely to discuss Wanting more ways to meet grad students and groups. Students who sought more structure meeting groups often talked about wanting ways to try different research, which aligns with the finding that these students also tended to enter graduate school with broader research interests and were less likely to have already connected with individual research groups. Additionally, we see that this subset of students also tended to have more struggles with Coursework and research in tension, as classes prevented them from dedicating time to their group search. This lack of time may have played a role in wanting more structured ways to meet research groups, as their schedules prevented them from doing so on their own time. Students who joined early occasionally cited this difficulty, but it was typically in regard to their desire to begin doing research in their lab rather than taking classes that they felt were not beneficial.
@@
Although many of the difficulties more commonly affected students who had not joined a group by their first semester, our results also reveal several drawbacks of joining a research group so early (the bottom of Table IV.5). Most notably, there was a higher prevalence of students being worried about Commitment to research before joining a group. Olivia, who planned to begin doing research with a professor as soon as she arrived in graduate school, feared that she could “get there and start talking about the research and doing it, and I might not like it at all… I don’t really know the solution for that.” Oftentimes students who had committed early to a group worried that they may be missing out on a better opportunity. This even included students who expressed high confidence in their choice of research. Ursula recalled that she “was trying not to pigeonhole myself too early still, even though I knew I loved this subfield.” In the end though, she committed to working with a professor whom she collaborated with during her undergraduate research.
@@
IV.6 Looking at the subset thinking about leaving
@@
As the first year progressed, both students experienced difficulties with Funding, Wanting more ways to meet grad students and groups, Coursework and research in tension, and Difficulty communicating with advisors, which contributed to their consideration to leave. Being denied from groups due to funding constraints negatively impacted sense of belonging for both students. Carmen was rejected from the first group they approached: “being my first attempt to reach out to a potential advisor, that was not a good feeling.” Brianna went to four professors in her department who all “said no, because they didn’t have funding for another grad student… personally, I thought I was cursed.” Carmen felt that the department did not do an adequate job facilitating interactions between graduate students, saying “I also felt like I lacked a sense of community… yeah, a lot of my concerns were community based.” Brianna recalled being confused as to “whether or not I had to find [a group] right away, as opposed to like, waiting out and focusing on trying to pass the courses first.” Opting to focus on coursework, she “was under the impression that could wait till my second year,” until the department informed her she needed a group by the summer, which made her feel “taken aback. I was scared.” Carmen described all of the first year responsibilities as “more overwhelming than I had expected.” Both described anxiety associated with seeking out help from faculty; as Brianna said, “I just didn’t know what was appropriate to ask and what not to ask… when you’re talking to people who hold the fate of your career in their hands, it’s like, I don’t know if they’re gonna think I’m dumb for asking this question. Because I’ve never experienced any of this before.”
@@
Leading up to graduate school, Tabitha and Kwan described experiences that were antipodal to those described by Brianna and Carmen. Both Tabitha and Kwan came into graduate school knowing exactly the kind of research they wanted to pursue, and had been in contact with prospective advisors before arriving to their programs. Kwan was guided to potential groups by his undergraduate research advisor, saying “some other candidate professors made my list but I screened them out by talking with my previous advisor. He was very supportive for applying to graduate school.” Meanwhile, Tabitha’s prospective advisor was a co-author on one of her undergraduate research papers: “So one of the professors at [this school] reached out to me and he was like, ‘Hey, I read your thesis. We would like to keep on working on this and make it an actual paper’ and I was like, sure.” Thus, when it came time for her to find a PhD program, this faculty member was a natural fit. She recalled him asking, “‘Have you to considered going to [this school]?’ Because he just happened to work here. Once again it’s like, you want me, you want to have me as your student? I’m not gonna say no.”
@@
Tabitha’s and Kwan’s research topics matched their interests excellently, but their work environments drove them to leave. From the outset Kwan felt that it was “inappropriate” for his advisor’s significant other to be working in the lab, and felt that the advisor treated him unfairly because of this dynamic. “It was a kind of big problem, a big burden. And one day we just yelled at each other, this kind of serious argument.” In his next meeting, his advisor said that their research style was “not fit” for him, and asked him to leave the lab. Kwan, an international student, recalled, “it was really stressful… I thought, why did I come to US?” He continued, saying “it was a kind very tough time for me. And I regretted about my decision to come to US a lot during my first year.” Meanwhile, Tabitha became disillusioned with her group due to the poor working environment fostered by the advisor. She said her advisor “created a very toxic work environment… always trying to gaslight me into believing I was not good enough to be here… [trying] to force me to work outside of my working hours.” Yet because she had come to the school specifically to work with this professor, she felt obligated to stay: “It took me a while to realize that I was not in a healthy work environment. Mostly because I was so, in a way, indebted to this professor for giving me the opportunity to come here. I think I thought, ‘Oh, yeah. He’s not great, but I’m here because of him.” More details of Tabitha’s case are detailed in our prior work [32]. Ultimately, support from her cohort helped to convince her that she would be able to succeed if she stayed: “I was doubting myself as a researcher. But my friends would tell me, no. It was you. You are doing the research.” Although not the focus of this study, this quote suggests how a student’s graduate cohort can also play an important role in supporting their sense of belonging.
@@
Other students in our study had joined groups early and subsequently switched, but only Tabitha and Kwan indicated that they considered whether they would be able to continue in their programs. Students who were motivated to switch groups in order to find research that better aligned with their interests did not indicate that their experience drove them to consider leaving their programs. Indeed, for students who came in with broad interests and joined a group in order to try out the research, switching was an expected and positive move. Consistent with prior literature on the impact of negative advising relationships, Tabitha’s and Kwan’s narratives demonstrate the outsized importance of group culture on student satisfaction. They also illustrate the negative influence that joining a group so early can have on students’ likelihood to stay in their programs if not properly supported. Tabitha said of her experience, “I definitely think that could have been avoided,” and she wished that that someone in the department had told her earlier to rethink working for her advisor. Reflecting on her search for a group, she emphasized to future students that “It’s a balance. It’s hard to balance having a good advisor and having a research field that you’re actually passionate about.”
@@
Figure 7: A model of characteristic pathways that physics PhD students take during the process of evaluating the congruence of a group’s research interests, exposing several underlying factors that contribute to students to thinking about leaving their programs. This model illustrates how these factors interact and influence a range of student outcomes. Student characteristics and outcomes on one side of the center dividing line are correlated with one other (e.g., broader interests are associated with searching during graduate school, which was correlated with thinking about leaving). Despite the left-right divide, we conceive of each variable on in the figure as existing on a spectrum. The grey text on the bottom left indicates an aspect of the model that we suspect exists, but was not present in our data. See Figure 8 for examples of how students in this study map onto the model.
@@ V Discussion
This study has examined the ways that physics PhD students’ research interests influence their search for a research group in graduate school. In this section, we synthesize our results in an explanatory model of student outcomes, shown in Figure 7. Grounded in data, the model highlights the connections between students’ backgrounds, their research interests, when they search for a group, and the consequences of searching for a group before or during graduate school. Moreover, by tracing students’ journeys back to their undergraduate years, this model offers unique insight into some of the ways that inequitable access to formative experiences can affect outcomes down the road in graduate school. Whereas prior research has correlated individual constructs (e.g., broad interests) with leaving [12], this model illustrates how these factors interact and influence a range of student outcomes. Student characteristics and outcomes on one side of the center dividing line are correlated with one other (e.g., broader interests are associated with searching during graduate school, which was correlated with thinking about leaving). Despite the left-right divide, we conceive of each variable on in the figure as existing on a spectrum. For instance, some students may have some access to undergraduate research that helps them develop somewhat stronger interests, but not enough to tell exactly what research they want to pursue.
@@
Figure 8 provides several examples of how students exemplified throughout the study are indeed well-described by diagram shown in Figure 7. Critically, our work offers insight into two characteristic pathways that we observed as placing students at a higher likelihood of leaving their programs, highlighted in Figs.8 a) and b). Figure 8a) represents the path exemplified by Brianna and Carmen. Both were first-generation students who entered graduate school unsure of what research they wanted to pursue. Thus their group search took place in graduate school, where they encountered a number of difficulties that impacted their overall sense of belonging in the program. Indeed, both considered leaving their programs before finding a research group. Meanwhile, Figure 8b) represents the path exemplified by Tabitha and Kwan, beginning with their strong research backgrounds that made them confident in the research they wanted to do in graduate school. Both students found a research advisor before they arrived in graduate school, primarily based on their research topic. However, upon working in their research groups during their first year, they recognized that their labs did not provide a working environment conducive to their success and considered leaving.
@@
Many other students’ experiences are captured in the model as well, as summarized in Figure 8. For example, Elena, Nina, and Alex are illustrative of the path shown in Figure 8c). Their undergraduate research played a defining role in determining the research they wanted to pursue in graduate school. And although this analysis primarily focused on the role of research interests in students’ search for a research group, many students in our sample were clearly cognizant of cultural factors in their choice of research group; for Elena, Nina, and Alex, their undergraduate institutions and research experiences gave them strong connections to their prospective groups. Elena and Nina had already worked on the same collaborations as their eventual graduate advisors, and Alex’s undergraduate advisor had previously worked with his prospective graduate advisor. Thus, this subset of students had insight into both the research and cultural aspects of their research groups early on in their search process, and they joined their labs with high confidence in having a good fit.
@@
Another notable pathway through the model is illustrated by Eric, Cole, and Kali in Figure 8d). We referenced Eric and Cole in Section IV.2 as having a broad set of interests coming into graduate school, but who joined a research group early anyway. Kali was also discussed in this section, indicating that she wanted to “explore” in her first year of graduate school. Thus, none of these students applied to their schools with the intention of working with a specific professor. Eric and Cole joined groups near the beginning of their graduate programs anyway, figuring that if it did not work out they would have time to switch. In effect, they were creating their own “rotation systems.” Meanwhile, Kali’s program already had an institutionalized rotation program, which allowed her to try a lab and switch. She credited the rotation system with helping her narrow her interests, saying “I don’t regret any of the rotations that I had done, because I know now that I don’t want to do that [type of research]… I wouldn’t have changed it.” We did not include Kali in our count of students who switched groups since it was a mandatory aspect of her graduate program, but we include her experience here since it most closely resembles the path shown in Figure 8d). Indeed, all three students found their first labs to be poor fits, predominantly due to the research. Therefore, they left and found new groups by the end of their first year summer.
@@
Lastly, in Figure 8e) we highlight several students who took a similar path as Brianna and Carmen but who never reported feeling isolation in their programs. Rather, they were able to navigate the search process in spite of the difficulties they experienced. Isaac and Ash had both some done undergraduate research but remained unsure of what exactly they wanted to do in graduate school. Both reported that having to balance coursework and research was a major issue during their first year and hindered their search, and both indicated that their departments did not provide much support in helping facilitate their group search. However, they never indicated that they were considering leaving their programs, and navigated their difficulties to find research groups by the end of the first year. One important factor that undoubtedly contributed to this difference in outcomes was Brianna and Carmen’s status as first-generation students. They cited this several times as a major reason for their struggle adjusting to graduate school overall. Difficulties navigating other aspects of graduate school while also worrying about finding a group undoubtedly exacerbated their negative feelings. We suspect that other factors contributed to the different outcomes as well, including departmental culture and the presence or absence of a strong graduate student cohort, but we did not explicitly investigate these elements in this work.
Figure 8: Several examples of students whose experiences searching for a group with congruent research interests that are well-described by the model in Figure 7. Each row represents a characteristic type of student experience, described alongside the student names in the study who fit that archetype. The paths listed under a) and b) describe students who our results suggested were at a higher risk of leaving their programs.
@@
Meanwhile, handbooks often have clear expectations for passing classes and qualifying exams during the first year. Among the schools in our sample, course requirements were the most clearly defined in all of the handbooks. The first-year graduate curriculum in physics is highly standardized, and the presence of classes in the first year is constant across physics graduate programs. Indeed, passing classes is often a core requirement for students to maintain good standing in their graduate programs. It is therefore unsurprising that students who come in relying on formal departmental guidance to navigate the first year may deliberately opt to delay looking for a group in favor of passing classes. However, this systematically disadvantages the 25% of physics graduate students who identify as first-generation college students, as this group is less likely to reach out for help on their own [93, 94]. Lacking guidance to the contrary, students may perceive experiences such as being denied access to a group due to funding seem like personal failings, contributing to feelings of isolation.
@@
Limitations and future research Interviewing students during their first and second years of physics graduate school allowed us to obtain deeper insights into their thoughts “in the moment” while looking for a group. However, in doing so we sacrificed some insight into these students’ outcomes later on in their graduate careers. The grey text in Figure 7 represents an area of the model that we were unable to investigate in this paper, but that we suspect to exist. We did not have enough data to investigate the outcomes for students who join a research group later on but experience a poor fit. Realizing that a lab is a poor fit can take time, as exemplified by Tabitha and Kwan who worked in their labs for over a year before switching. Students who join a group later in their first year might not realize that the fit is not great until late into their second year. By then, the cost of switching might feel too great, and could leave them feeling as though they have no choice but to stay in the poor research environment or leave the program. More research into these students’ experiences would be needed to understand how they fit into our model.
@@
The goal of this paper from the outset was to emphasize the role of students’ research interests in how they find a group. Thus, a clear limitation of the results presented here is that they do not offer as much insight into how students evaluate group cultures and working environments. A positive interpersonal dynamic in a research group is clearly essential for students to thrive, but how students weigh these considerations in comparison to their research interests remains unexplored within our data. Previous work in the context of biology graduate education indicates that students weighed both factors heavily when choosing a group [28], and we see indications of that in our data as well. However, a fuller description is needed to understand how students approach finding a group with a positive culture.
@@
Lastly, our data does not include students who left their graduate programs. Their perspectives on the physics graduate experience are critically important, but are often difficult to collect due to poor tracking of students once they leave their programs [12]. Although our data included students who reported a strong consideration of leaving, they ultimately stayed. Gaining a more accurate characterization of how finding a research group impacts student retention, satisfaction, and leaving will require more work to gather data from those who did not complete the PhD. It is therefore important to recognize that the trends observed in this study may not adequately represent these students [99]. Future research should also attend to how it is experienced across demographic groups and within different institutional contexts, as our study was limited in these regards. Our sample was diverse in many ways, but our protocol did not explicitly probe how students felt their identities impacted their group search. Moreover, this analysis did not systematically examine differences across race, gender, or international status regarding time to join groups and development of research interests. And although first-generation status came up naturally during our study played a significant role in several students’ narratives, this was not our central focus. Future research must attend to how physics students’ identities shape the first year of the doctoral experience in order to provide more holistic and equitable support for all students.